# Automated Modeling Method for Pathloss Model Discovery

### v3

@@ VII Conclusion
This work also revealed several challenges to be addressed in the future. More specifically, utilizing a clear form of path loss, the non-logarithmic form, may enable more precise modeling. The performance-interpretability tradeoffs with KANs could likely be improved by enhanced symbolic mapping. Further methodological tools to assess the validity of the composition graph learned by the KAN may ease the adoption of the proposed method. Finally, extending both the KANs and DSR methods to cover diverse scenarios will be essential to validate their robustness and adaptability. These advances have the potential to contribute significantly to the development of intelligent, adaptive modeling frameworks in the field of wireless communication systems.
Acknowledgements
This work was supported by the Slovenian Research Agency (P2-0016) and the European Commission NANCY project (No. 101096456).

### v4

@@ VII Conclusion
This work also revealed several challenges to be addressed in the future. More specifically, utilizing a clear form of PL, the non-logarithmic form, may enable more precise modeling. The performance-interpretability tradeoffs with KANs could likely be improved by enhanced symbolic mapping. Further methodological tools to assess the validity of the composition graph learned by the KAN may ease the adoption of the proposed method. Finally, extending both the KANs and DSR methods to cover diverse scenarios will be essential to validate their robustness and adaptability. These advances have the potential to contribute significantly to the development of intelligent, adaptive modeling frameworks in the field of wireless communication systems.
Acknowledgments This work was supported by the Slovenian Research Agency (P2-0016) and the European Commission NANCY project (No. 101096456).